OpenStudy (anonymous):

IS MY ANSWER RIGHT? answer d If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable y is eliminated in the first and second equations, then the second and third equations. 3x - y + 3z = 1 4x + y - 6z = 3 4x - 2y + 3z = -3 (4 points) Answer 7x - 3z = 4 12x - 9z = 3 7x - 3z = 4 4x - 6z = 3 3x + 3z = 1 12x - 9z = 3 3x + 3z = 1 4x - 6z = 3

OpenStudy (anonymous):

A is the correct option.. Hope this helps

OpenStudy (anonymous):

add eqs 1 & 2 then u get 7x - 3z = 4 multiply eq-2 by 2 and then add to eq-3, as below 2*(4x + y - 6z) = 2*3 => 8x + 2y - 12z = 6, now add to eq-3, then u get 12x - 9z = 3 so, the correct answer is A option..

OpenStudy (anonymous):

i love you thank you thank you